OpenStudy (anonymous):

a^3=5/2 got help and they said it was cuberoot20/2 but i don't get why.

OpenStudy (anonymous):

What is the exact question that u have??

OpenStudy (anonymous):

2a^6-3a^3-5=0

OpenStudy (anonymous):

let u=a^3, 2u^2-3u-5=0 (2u-5)(u+1)=0 u=5/2 or u=-1 which leads to a^3=5/2 or a^3=-1 a=cuberoot(5/2) or a=-1

OpenStudy (anonymous):

thanks! i got those answers but the answer sheet said cuberoot(20)/2 so i wasn't sure. thanks for answering!

OpenStudy (anonymous):

welcome :)